### v1

@@ 5 SRON Netherlands Institute for Space Research, Niels Bohrweg 4, 2333 CA Leiden, The Netherlands
AMKID - a large KID-based camera at the APEX telescope
The thermal emission at sub-millimeter wavelengths carries unique information in many astronomical applications ranging from disks and planet formation around young stars, to galaxy evolution studies at cosmological distances. Advancing on the mapping speed to detect this faint emission in ground-based astronomy has been a technical challenge for decades.
The APEX Microwave Kinetic Inductance Detector (AMKID) camera was designed to accomplish this task. The instrument is a wide field-of-view camera based on kinetic inductance detectors. It is installed on the 12 meter APEX telescope in Chile at 5.100 meters above see level. The instrument operates dual color, covering simultaneously the 350 GHz and 850 GHz atmospheric windows. It has a large field-of-view of 15.3’x15.3’, and an unprecedented number of pixels: 13.952 detectors in the high frequency band and 3.520 detectors in the low frequency band. Here we present a complete description of the instrument design and construction together with results of the successful low frequency array (LFA) commissioning campaign executed during the last year.
The LFA performance is in good agreement with design parameters, with detector sensitivity of 2.2 mK and diffraction limited beam sizes of 17.0”. On-sky measurements demonstrate a sensitivity of 70-90 mJy per detector when operating under good atmospheric conditions (PWV below 1.0mm). With this performance the LFA regularly achieve a mapping sensitivity of 25 mJy when mapping a square degree in an hour.
AMKID on APEX with its dual color observing capabilities, high sensitivity, large field-of-view and high angular resolution holds the promise to open a new range of science with the APEX telescope.
@@
Until its de-commissioning in 2020 the wide-field direct detection camera LABOCA (Kreysa_2003; Siringo2009), operating up to 295 horn-coupled thermistors in the 870µm atmospheric window, has been the most successful instrument at the APEX telescope, described in (Guesten_2006), leaving a fantastic legacy with close to 500 scientific publications. LABOCA’s noise-weighted point-source sensitivity per channel of 55 mJy, after removal of correlated sky noise, served as gold standard for subsequent developments, but the underlying technology also limited the number of detectors in the camera’s field of view.
A path to more efficient large-scale observations, towards focal-plane arrays with several kilo-pixels at affordable costs, did open in the early 2000s with the introduction of the kinetic inductance detector (KID); (Day_2003). Since then, a number of KID-based cameras have been deployed for operation in the mm regime, such as NIKA and NIKA2 at the IRAM-30m telescope (Monfardini_2010; NIKA2), or MUSIC at the Caltech Submillimeter Observatory (Maloney_2010). Other instruments, also covering the submillimeter regime, are currently at different stages of development as the TolTec (Wilson_2020), BLAST-TNG (Lourie_2018) or P-CAM (Vaskuri_2025).
Building on these promises MPIfR with SRON have launched an ambitious project to develop and operate a wide-field, KID-based submm-camera for the APEX telescope: the APEX microwave KID camera (AMKID), initially described in Esteras_2015. It is a dual-color instrument, operating in the 870 and 350 m atmospheric window simultaneously, thereby taking advantage of the superb atmospheric conditions at Llano de Chajnantor and the excellent sub-millimeter performance of the telescope. The detectors are lens coupled MKID devices, filling the 15 arcmin field-of-view of the telescope in technically densest possible arrangement, meaning several thousands of pixels detectors per array. Taking advantage of the site, operation of the detectors is close to be background limited.
In this article we report on the overall instrument design and focus on the successful commissioning of the low-frequency camera. The high-frequency complement will follow shortly, as will the opportunity for polarization observations.
For a more detailed discussion on the scientific motivations we refer the reader to the extensive literature on this subjects, which range from disks and planet formation around young stars (e.g. Andrews_2018; Holland_2017), to large scale Milky Way studies (e.g. Schuller_2009; Molinari_2010), or galaxy evolution at cosmological distances (e.g. Weiss_2009; Oliver_2012) just to mention some relevant examples. The promises of mapping large scale structures on southern skies, be it galactic or cosmological, with the powerful state-of-the-art AMKID arrays installed at the APEX on one of the world’s outstanding submm-sites are unique and exciting.
@@ 2 Instrument Overview
The AMKID instrument is a wide-field sub-millimeter camera. It operates at 870 m (345 GHz) and 350 m (850 GHz) using its Low Frequency Array (LFA) and High Frequency Array (HFA) respectively. Both arrays are diplexed in polarization by a free-standing wire polarizer and can be used simultaneously. The instrument is designed for operations at the 12 meter APEX telescope, making use of the superb observation conditions at the Chajnantor plateu. Precipitable water vapour (PWV) below 0.5 mm is observed over several weeks a year, allowing regular operation in the 350 m band. In addition the telescope surface quality is regularly monitored and adjusted via holography, keeping it in the range of 15-20 m RMS over the complete observation season.
The instrument optics is designed to make use of the complete field-of-view avalaible at the telescope (15.3’ x 15.3’). The field is sampled with an unprecedented number of pixels: 3,520 pixels in the low frequency band and 13,952 pixels in the high frequency band, using a pixel spacing of 1.2/D = 18” at the LFA and 1.5/D = 9” at the HFA. The instrument optics is near the diffraction limit, with nominal half maximum beam sizes of 17.0” at the LFA and 7.5” at the HFA.
@@ 3 Detector technology
KIDs were pioneered in the early 2000s by the group of J. Zmuidzinas at Caltech (see Day_2003). They are superconducting microwave resonators designed to efficiently absorb radiation exceeding their gap frequency, which is 90 GHz for Aluminum. Upon radiation absorption Cooper pairs are broken into single particle excitations called quasiparticles, resulting in an increase in both the (kinetic) inductance and the extremely small resistance of the superconducting film at microwave frequencies. As a result, the resonance frequency decreases which can be read-out very accurately using a single microwave probe tone at the unperturbed resonant frequency.
@@
To eliminate thermal excitations and excess noise KIDs need to be operated far below the superconductor critical temperature, K for Aluminum. At these temperatures all charge transport is achieved by Cooper pairs, which have zero resistance, but a finite (kinetic) inductance associated with the electron mass.
@@
The main advantage of this technology is that individual KIDs can be designed to each have a specific resonance frequency, so that each detector in a KID array is associated with a single resonance dip, as shown in Fig. 3. As is clear from this figure many KIDs can be multiplexed in frequency space and read using a single microwave line, a single cryogenic amplifier, and a single room-temperature back-end. This concept
allows to build large arrays, opposed to other previous technologies where each detector
needed its own wiring, limiting the number of pixels.
The AMKID detectors are NbTiN-Aluminum lens-antenna coupled hybrid KIDs as shown in Fig. 2. These detectors, fabricated on high resistivity silicon, are described in Jansen_2013 for the LFA and in Ferrari_2018a for the HFA. They reach photon noise limited performance for the loading powers relevant for AMKID combined with a high coupling efficiency. The Appendix A offers a more detailed overview on the used detector technology.
To allow for small detector sizes we read-out the KIDs in the frequency band of 4 to 8 GHz. Resonances are designed to operate with depths between 3 and 6 dB and Q factor of the order of 2.1E4 under the
@@
target optical load, as seen in Fig. 3. The monolithic silicon lens arrays are fabricated using laser ablation, with a typical fabrication accuracy of 5 m in form factor. Each lens array is coated with a anti-reflection coating made from Parylene-C, as described in Ji2000. The detector efficiency is estimated using numeric simulations to be 84% and 76% at the low and high frequency bands. Between the lens and the planar antenna we use a capacitive grid which absorbs unwanted in-chip stray light (see Yates_2017).

### v2

@@ 5 SRON Netherlands Institute for Space Research, Niels Bohrweg 4, 2333 CA Leiden, The Netherlands
AMKID: A large KID-based camera at the APEX telescope
Thermal emission at submillimeter wavelengths carries unique information for many astronomical applications, ranging from disks and planet formation around young stars to galaxy evolution studies at cosmological distances. Advancing the mapping speed to detect this faint emission in ground-based astronomy has been a technical challenge for decades.
The APEX Microwave Kinetic Inductance Detector (AMKID) camera was designed to accomplish this task. The instrument is a wide field-of-view camera based on kinetic inductance detectors. It is installed on the 12-meter APEX telescope in Chile at 5,100 meters above sea level. The instrument operates in dual color, covering the 350 GHz and 850 GHz atmospheric windows simultaneously. It has a large field of view of 15.3’ x 15.3’ and an unprecedented number of pixels: 13,952 detectors in the high-frequency band and 3,520 detectors in the low-frequency band. Here we present a complete description of the instrument design and construction, together with results from the successful low-frequency-array (LFA) commissioning campaign executed over the last year.
The LFA performance is in good agreement with design parameters, with detector sensitivity of 2.2 mK and diffraction-limited beam sizes of 17.0”. On-sky measurements demonstrate a sensitivity of 70-90 mJy per detector under good atmospheric conditions (PWV below 1.0mm). With this performance the LFA regularly achieves a mapping sensitivity of 25 mJy when mapping a square degree in one hour.
AMKID on APEX with its dual-color observing capabilities, high sensitivity, large field of view, and high angular resolution promises to open new scientific opportunities with the APEX telescope.
@@
Until its decommissioning in 2020, the wide-field direct-detection camera LABOCA (Kreysa et al. 2003; Siringo et al. 2009), operating up to 295 horn-coupled thermistors in the 870µm atmospheric window, has been the most successful instrument at the APEX telescope (Güsten et al. 2006), leaving a fantastic legacy of nearly 500 scientific publications. LABOCA’s noise-weighted point-source sensitivity per channel of 55 mJy, after removal of correlated sky noise, served as the gold standard for subsequent developments, but the underlying technology also limited the number of detectors in the camera’s field of view.
A path to more efficient large-scale observations – toward focal-plane arrays with several kilopixels at affordable costs – did open in the early 2000s with the introduction of the kinetic inductance detector (KID); (Day et al. 2003). Since then, a number of KID-based cameras have been deployed for operation in the millimeter regime, such as NIKA and NIKA2 at the IRAM-30m telescope, and MUSIC at the Caltech Submillimeter Observatory (Monfardini et al. 2010; Adam et al. 2018; Maloney et al. 2010). Other instruments covering the submillimeter regime – such as the TolTec, BLAST-TNG, and P-CAM (Wilson et al. 2020; Lourie et al. 2018; Vaskuri et al. 2025) – are currently at different stages of development.
Building on these promises, MPIfR with SRON have launched an ambitious project to develop and operate a wide-field, KID-based submillimeter camera for the APEX telescope: the APEX microwave KID camera (AMKID), initially described in Esteras (2015). It is a dual-color instrument, operating simultaneously in the 870 and 350 m atmospheric windows, thereby taking advantage of the superb atmospheric conditions at Llano de Chajnantor and the telescope’s excellent submillimeter performance. The detectors are lens-coupled MKID devices, arranged in the densest technically possible configuration to fill the telescope’s 15.3 arcmin square field of view, with several thousands of pixels detectors per array.
In this article we report on the overall instrument design and focus on the successful commissioning of the low-frequency camera. The high-frequency complement will follow shortly, as will polarization observations.
For a more detailed discussion of the scientific motivations, we refer the reader to the extensive literature on this subject, which spans disks and planet formation around young stars (e.g., Andrews et al. 2018; Holland et al. 2017), large-scale Milky Way studies (e.g., Schuller et al. 2009; Molinari et al. 2010), and galaxy evolution at cosmological distances (e.g., Weiß et al. 2009; Oliver et al. 2012), among other relevant examples. The promises of mapping large-scale structures on southern skies, be it galactic or cosmological, with the powerful state-of-the-art AMKID arrays installed at the APEX on one of the world’s outstanding submillimeter sites are unique and exciting.
@@ 2 Instrument overview
The AMKID instrument is a wide-field submillimeter camera. It operates at 870 m (345 GHz) and 350 m (850 GHz) using its Low Frequency Array (LFA) and High Frequency Array (HFA) respectively. Both arrays are diplexed in polarization by a free-standing wire polarizer and can be used simultaneously. The instrument is designed for operations at the 12-meter APEX telescope, making use of the superb observation conditions at the Chajnantor plateau. Precipitable water vapor (PWV) below 0.5 mm is observed over several weeks per year, allowing regular operation in the 350 m band. In addition, the telescope’s surface quality is regularly monitored and adjusted via holography, maintaining 15-20 m RMS over the entire observation season.
The instrument optics is designed to make use of the maximum square field of view available at the telescope (15.3’ x 15.3’). The field is sampled with an unprecedented number of pixels: 3,520 pixels in the low-frequency band and 13,952 pixels in the high-frequency band, using a pixel spacing of 1.2/D = 18” at the LFA and 1.5/D = 9” at the HFA. The instrument optics is near the diffraction limit, with nominal half maximum beam sizes of 17.0” at the LFA and 7.5” at the HFA.
@@ 3 Detector technology
KIDs were pioneered in the early 2000s by the group of J. Zmuidzinas at Caltech (see Day et al. 2003). They are superconducting microwave resonators designed to efficiently absorb radiation exceeding their gap frequency (90 GHz for aluminum). Upon radiation absorption Cooper pairs are broken into single particle excitations called quasiparticles, resulting in an increase in both the (kinetic) inductance and the extremely small resistance of the superconducting film at microwave frequencies. As a result, the resonance frequency decreases, which can be read out very accurately using a single microwave probe tone at the unperturbed resonant frequency.
@@
To eliminate thermal excitations and excess noise, KIDs must be operated far below the superconductor critical temperature ( K for aluminum). At these temperatures all charge transport is achieved by Cooper pairs, which have zero resistance but finite (kinetic) inductance associated with their electron mass.
@@
The main advantage of this technology is that individual KIDs can be designed with specific resonance frequencies, so that each detector in a KID array corresponds to a single resonance dip, as shown in Fig. 3. As shown, many KIDs can be multiplexed in frequency space and read using a single microwave line, a single cryogenic amplifier, and a single room-temperature backend. This concept allows us to build large arrays, contrary to previous technologies where each detector
required individual wiring that limited the number of pixels.
4 KID detectors fabrication
The AMKID detectors are NbTiN-aluminum lens-antenna-coupled hybrid KIDs, as described in Janssen et al. (2013) for the LFA and in Ferrari et al. (2018b) for the HFA. The detector geometry is shown in Fig. 2, with dimensions specified in Table 1. They are fabricated on a high-resistivity silicon wafer, as described in Thoen et al. (2017) and Bos et al. (2017). Both high- and low-frequency detectors are very similar, with the main difference being their antenna dimensions, which scale with operating frequency. This section gives some further details about the detectors’ design and fabrication.
The coplanar waveguide (CPW) resonators are made of a 400 nm film of NbTiN with a K. At the open end, the resonator is coupled to the readout line; here, the CPW line is widened to reduce 1/f noise (see Gao et al. 2008). At the shorted end of the resonator, a section of the CPW line narrows significantly, with the central line made of aluminum. The length of this section is 30% of the KID’s length (i.e., about 1-2 mm long, depending on the KID resonant frequency). This scaling improves the readout frequency scatter of the KIDs, since any variations in aluminum line width will have the same fractional effect on all detectors. Another advantage is the frequency-independent responsivity and photon noise level.
The resonator is coupled to a twin-slot antenna (see Zmuidzinas and LeDuc 1992; Filipovic et al. 1993), which couples the radiation from the telescope into the NbTiN-aluminum hybrid CPW by means of a silicon lens mounted on the chip’s backside (see Fig.2b). Due to the antenna coupling scheme, AMKID detectors are single-linear-polarization detectors. Quasiparticle creation only occurs in the aluminum line, since the gap frequency of NbTiN is 1.1 THz and thus provides completely loss-free material for the resonator and antenna.
@@ 4 KID detectors fabrication
The monolithic silicon lens arrays are fabricated using laser ablation, with a typical fabrication accuracy of 5 m in form factor. Each lens array is coated with a antireflection coating made from Parylene-C, as described in Ji et al. (2000). Numeric simulations restimate detector efficiency at 84% and 76% in the low- and high-frequency bands. Between the lens and the planar antenna, we place a capacitive grid that absorbs unwanted in-chip stray light (see Yates et al. 2017).